OpenStudy (anonymous):

The National Association for Women in Science asked recent high school grads if they had taken certain science classes. Of those surveyed, 45 said they had taken physics, 67 said they had taken chemistry and 22 said they had taken both. Seventeen said they had taken neither. How many recent high school grads were surveyed? please Show your work.

OpenStudy (lalaly):

17 + 22 + [ 45-22] + [ 67 - 22]

OpenStudy (anonymous):

what is the answer to slove this!

OpenStudy (anonymous):

yes i think shes right 17+22+(45-22)+(67-22) '45 said they had taken physics, 67 said they had taken chemistry and 22 said they had taken both' so you should subtract 22 from physics and chemistry

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

trying to figur out howto slove this !

OpenStudy (lalaly):

do u need help?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

17+22+(45-22)+(67-22) ok whats 45-22 and 67-22?

OpenStudy (anonymous):

45-22=23 and 67-22=45

OpenStudy (anonymous):

so 17+22+23+45=?

OpenStudy (anonymous):

107

OpenStudy (anonymous):

so the answer=107